What is molar polarization? What types of polarizations contribute to total molar polarization (Pm)? Give an equation for total molar polarization.
1
Expert's answer
2019-11-11T07:47:25-0500

Molecules whether polar or non polar are electrical in nature due to the presence of nuclei and electrons. LThey are neutral because their net charge is zero, the number of nuclear charges being equal to the number of electronic charges. Inthe case of non polar molecules, the centers of positive and negative charges coincide. However, when such a molecule is 

placed in an electric field, there is a displacement of electric centers. The center of positive charges shifts towards one end while that of negative charges shifts towards the other end of the molecule. As a result, a dipole is induced in the molecule. 

This is called distortion polarization. Polar molecules have a permanent dipole. Such molecules are said to have orientation 

polarization.

Thus total polarization „P‟ in such case is given by 


"P=P_o+P_1"

Where "P_o\ and\ P_1" represent orientation & induced polarization.



"P_m=\frac{D-1}{D+2}\times\frac{M}{\rho}"